OpenStudy (anonymous):

simplify x^3/2(x+x^5/2-x^2)

OpenStudy (anonymous):

multiply happily along adding the exponents as ye go

OpenStudy (anonymous):

=x^5/2 + x^4 - x^7/2

OpenStudy (anonymous):

\[x^{1+\frac{3}{2}}+x^{\frac{3}{2}+\frac{5}{2}}-x^{\frac{3}{2}+2}\]

OpenStudy (anonymous):

get what jimmyrep said